OpenStudy (mysticfury):

can someone help me figure out how to graph this equation -> 2x − 5y = −15

OpenStudy (anonymous):

I was just about to tag you @563blackghost :P XD

OpenStudy (smartnerd1111):

look at the pic

563blackghost (563blackghost):

You have to get y by itself...so we would first subtract 2x to get it to the other side... \(\LARGE{2x-5y=-15 ~\rightarrow ~-5y=-2x-15}\) Next we would divide by -5 to get y by itself.... \(\Huge{\frac{-5y}{-5}=\frac{-2x-15}{-5}}\) \(\Huge{~~~~~~~~\downarrow}\) \(\Huge\color{red}{y=\frac{2}{5}x+3}\)

OpenStudy (smartnerd1111):

we got the same answer yay

OpenStudy (smartnerd1111):

we got the same answer yay

563blackghost (563blackghost):

^^

OpenStudy (mysticfury):

alright i gotcha now how do i graph the slope intercept form? i need to figure out if it has a negative or positive rate of change, or do i even need to graph it?

563blackghost (563blackghost):

Well the slope does not contain a negative so it would be positive... To graph you must know the format of slope-intercept form... \(\Huge{y=mx+b}\) \(\Huge{~~~~~~slope~~y-intercept}\) m=slope b= y-intercept So first graph the y-intercept then place the slope...

OpenStudy (mysticfury):

okay so if the slope would be -3/4 thats negative rate of change and if it dosent have the - its positive? so this would be negative then? y = −4x + 9

563blackghost (563blackghost):

Correct ^^

OpenStudy (mysticfury):

alright thank you very very much!

563blackghost (563blackghost):

np :)